OpenStudy (breanna.rae):

PLEASE HELP write an equation in point slope form of the line through point J(4,1) with slope -4

OpenStudy (steve816):



OpenStudy (breanna.rae):

That the answer?

OpenStudy (ryandean123):

y- y1 = m(x - x1) given m and (x1, y1) y - 1 = -4 ( x- 4)

OpenStudy (breanna.rae):

so its y - 1 = -4 (x-4)? I'm not good at this at all I'm sorry

OpenStudy (steve816):

Yup that is the correct answer! ^^

OpenStudy (breanna.rae):

thank you both!

OpenStudy (ryandean123):

HAHHAHA I WIn im smart

OpenStudy (steve816):

No problem!